OpenStudy (anonymous):

if a drug comes in a solution of 0.125 g / mL, what volume will you measure for a dose of 0.25g?

OpenStudy (whpalmer4):

\[\text{dose}=\text{concentration}*\text{volume}\] your dose \(= 0.25\text { g}\) your concentration \(= 0.125\text{ g/mL}\) volume = ______ mL

jhonyy9 (jhonyy9):

just a little help than you see 0,25 g is exactly double of 0,125 what is the concentration so this mean that volume equal = ?

jhonyy9 (jhonyy9):

do you understand it ?

jhonyy9 (jhonyy9):

than 0,125 g in 1 ml so 0,25 g -- in ? ml

jhonyy9 (jhonyy9):

are you here ?

jhonyy9 (jhonyy9):

so than you multiplie by 2 *0,125 g = 0,250 g so than just multiplie this 1 ml volume too by 2 and will get the necessary volume for this 0,250 g drog - right ?

OpenStudy (whpalmer4):

\[\text{dose}=\text{concentration}*\text{volume}\]or to save a bit of typing\[d = c*v\]You want to find volume, so we can rearrange that to be \[\frac{d}{c} = v\]by dividing both sides of the equation by \(c\) \[v = \frac{0.25\text{ g}}{0.125\text{ g/mL}} = \frac{0.25\cancel{\text{ g}}}{0.125\cancel{\text{ g}}\text{/mL}} = \frac{0.25}{0.125}\text{ mL}\] I trust you can do the numerical part of the division.